Commander Report November 8th

A great challenge for me is to prepare the meals in the first days of our simulation. In the first days, I considered had did delicious meals. I never had prepare a meal as I had done here. So, for me, it is a great learning that I will bring back to Brazil. Thank you MDRS and Mars Society for makes emerges in me some new competencies applied to my life. J

7h45m – I spent 1 hour walking in the MDRS corridors, in a distance of 6,4 km or 4 miles.

8h55m – In Hanksville is doing 1oC. or 34oF.

9h – Preparation of the breakfast.

10h30m – We still in breakfast. One of the main topics of our debates is related to internet. Just in the first hours, our internet just consumed 290 MB from 550 MB available. So, We are very worried about that. K

10h45m – My soundtrack today is Megadeth album Symphony of Destruction during the time when I’m writing the book related to Mars and Sustainability.

12h – 13h – Shannon Rupert (MDRS Director) presented to us a very detailed instruction about the use of the All-Terrain Vehicles (ATV). Shannon is always very friendly with us.

13h – Start to prepare the lunch.

13h20m – Stop to help in the EVA preparation.

13h30m – Atila, Brandon and Carmen develop an EVA activity using the ATV. The returned 15h30m.

13h40m – Return to prepare the lunch.

14h40 – Just preparing the lunch.

15h50m-16h50m – Lunch. Our meal today was brown beans, rice with broccoli and beef dehydrated.

17h20m – I just walked more 30 minutes of 4,8 km or 3 miles, completing 11,2 km of walking, or 7 miles.

We all considered that the day was so short today and the time pass so fast.

Sol Summary November 8th

MDRS Daily Summary Report for sol 5

Summary Title: Time is running fast today.

Mission Status: The day was very busy today. All the crew members had something to do.

Sol Activity Summary: Prepare the meal today was so painful. I never cooked beans. It looks like old. I had some problem with it. Our crew had a good time today. All doing different and complementary activities. We are all well motivated and keeping a good relationship. Thank You for all presenting the wishes relate with us.

Look Ahead Plan: Tomorrow, We are waiting the Brazilian reporters of TV Globo at 9 AM. We are planning an EVA during their time here. We are just prepared to receive them.

Anomalies in work: Nope.

Weather: 6 C, 43 F

Crew Physical Status: Fine.

EVA: today happened EVA with the ATV.

Reports to be file: We sent in CapComm.

Support Requested: None.

[end]

Science Report November 8th

Mars today has been really great, with bright sunshine and now a cozy hab to sleep in!

Things in the lab have been really great! The germination tests of the raddish seeds are going great. 3/5 in one plate and 3/5 in the second plate have presented a positive response, so the seeds are viable. Raddish seeds are also really fast at germinating, but this new brand we bought is particullarly fast. This will be great for the lab tests.

The new autoclave is now at the lab, and we have plenty of destillated water, so that will help a lot with the different biological experiments. The new microscope is also set up and ready to go! It also incorporates a new cam for our sights in it, so I would like to ask where should I install the drive to use the cam. I fear that this drive is just a one-time software, so if I install it in my computer, it would be the only time it could be installed. That happens quite a lot with other copyrighted softwares. I tried to install it in the hab laptop, but we couldn’t find the cable to connect it.

At today’s EVA, our crew as able to recollect more soil for their experiments. A crewmate is trying to tipify the soil, while other project involves the germination of different seeds in two types of soil.

There’s also this idea some of us had of using the soapy water for some seeds. I’m running a pilot test with raddish seeds. It would be interesting if we find a way of making that dirty soapy water from the dishes useful!

I’ll attach some photos of the experiments in the lab today. The first two ones are from my raddish experiments. The last ones are from the pilot experiment with the dirty water from the dishes.

We also tried to organize the lab better. There are now labels in each drawer and we are trying to set places for the equipment. There is a table for running microbiological experiments and another for geological ones. We are hoping to make the lab a great and ordered space for future crews!




Operations Report November 8th

SOL: 5

Carmen Atauconcha

Non-nominal systems: Working

Generator: On

Solar: For now, not given power to the Hab,

Diesel : 64%

Propane – 62psi

Gasoline (5 Gallon containers for ATV) – 0 gallons

Water (trailer) – empty

Water (static) –

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 519 gallons

ATV’s Used: Yes

Oil Checked: motos y carros: Yes

Atv # Fuel Used Gals: 2 gallons

Tires Status: Perfect

# Hours Used Day: 4hours

Phobos used: No

Hours:0

Beginning charge: –

Ending charge: –

Currently charging: –

Deimos used: Yes

Hours: 4hrs

Beginning charge: 4pm

Ending charge: Continuing charging

Currently charging: Yes

HabCar used?: No

Notes and Comments:
Last night was a little cold, we are having problem with the heater., because when we turn it on, it heats, but few hours later the air is cold again. So, we have to turn it off and then turn it on again to keep us warm. Do you know a solution for this problem?

At the same time, we did an EVA session today. The blue moto is not working. In that way, we did the EVA session with the deimo and the red moto. We do not had any problems with these vehicles. In contrast, we had problems with the radios. All radios are working, however, approximately 2.5 km far from the Hab, we could not had any communication with the Hab. We do not know why that is. So, please send us the ratio of distance in which we could keep the communication.

Tomorrow we are going to use the solar.

EVA Report November 8th

We started EVA at 1:35 PM, instead of the planned 1:00 PM. The delay was due to an extended period for our ATV training to take place. The EVA crew included myself, Carmen, and Atila. Drove north and spotted vegetation growth along road and marked the location. We tried marking the location with a GPS, but we found it was impossible to read off the screen with the helmets on, since the screen was too scratched up. Instead, I determined the coordinates by closely examining the planning map after the EVA. We Took several pictures of the area along with the surrounding terrain

We continued north and passed Tank Wash due to navigational Error. We ended up taking the road all the way to the Quarry Site. We retraced our steps southward until we found the reservoir dam, which was the most distinguishable feature I could use to place our location.

-Stopped at coordinates:

4252820

518590

We then slowly drove northward, until we found the correct wash

-Stopped at coordinates:

4253440

518430

We made our way to a talus slope that contained broken down rock from the cliff above us. Upon digging through the cemented layers, we found three distinct layers within only 20 cm below the surface.

– Coordinates of Sample Location #3:

4253680

518600

There was a white-colored layer in the first 7 cm, then a green layer below it that was 3-4 cm thick. Below that was a dark brown / black soil lightly peppered with white grains. We collected samples from each of these layers and photographed the area and our activities. We loaded up our samples and returned to the hab. Before we entered the airlock, Carmen checked the Diesel and Propane tank levels. We began recompression in the airlock by 3:45 PM.

For the next EVA, we need to improve our radio Comms. I’ll instruct the crew members on words and phrases to use on the radio, such as “affirmative”, “negative”, “over”, “loud and clear”, etc. Also going over the phonetic alphabet may come in handy. In order to reduce the chance of another navigational error, I’ll have to take very special notice of the surface features surrounding my intended destination before we depart. I’d also like to request obtaining another GPS device that we can read on the field so we can record our sample locations in the future.

Green Hab Report November 8th

Green Hab Officer Atila Meszaros

Sol 5 – 08/11/2017

Ambient (no heating or/cooling): none

Ambient with window/door open: door opened at 10:53; closed at 16: 20

Shade cloth on/off: off

Average temperature:

Time 10:53 12:00 13:40 18:22
Temperature inside (°C) 33 34 34 12
Temperature outside (°C) 5 5
% Humidity 16 16 16 16

*Time when seeds of a former experiment where watered. (There are tomatoes, melons, peppers, cucumbers, and beans)

Changes to crops: None

Daily water usage for crops: 2 ½ galons

Changes to research plants: none

Daily watering and amount of water used: 864 mL approximately

Sol 5 was a pretty lazy day at the Green Hab (GH). Yesterday I worked till 22:00 at the GH and then I was looking for some guides to build a homemade Drip Irrigation System, so I woke up pretty late, plus we had the training for the Rovers and the ATV’s and a 3-hour EVA. It all went awesome! But I didn’t have time to make the repairs I wanted to make on the GH. However, I did repair the GH door that I discovered was a bit lose.

During the last hours of Sol 4, I finished the preparations for my project. Half of the seeds remain at the GH; the other half is in the growing chamber that’s on the Lab. For all the experiment I’m using less than a liter of distillated water per day.

The project wants to prove that Peruvian native crops can, indeed, germinate and grow on Mars soil. I’m working with quinoa, amaranthus and mustard (as a control).

I’ll attach you a picture of the pots!

As I’m aiming every Sol, tomorrow I’m going to rearrange all the GH, because it’s a mess. I’ll let you know how that worked.

Support/supplies needed: none

Sol Summary November 3rd

Sol Summary Report for sol 0.3

November 3, 2017

And success!

Shannon Rupert, MDRS Director

Mission Status: All work party objectives were addressed. See Kay’s
engineering report for details.

Sol Activity Summary: Another busy day for us. We worked on the
objectives we set for ourselves. Highlight of my day was getting the
generator running. It had been offline since June. Kevin and John
spent much time today in the GreenHab and so I am happy to report the first crew will have seeds to tend. Lowlight of my day was the resignation of DG Lusko, who has been our invaluable local support
here in Hanksville for about a decade. For me personally, he has been
the person I have relied upon for so much since I became director of
the station. I have also learned a lot from him and after we got the
generator running through my plan, I realized I can actually fly after
being kicked out of the nest.

Look Ahead Plan: Crew 182 arrives tomorrow at noon! We have a lot of
clean up to do tomorrow! John L. will be leaving us in the morning,
Kevin in the afternoon. Kay and John S will stay until Sunday and
overnight with the new crew.

Anomalies in work: There are a lot of missing/misplaced items here at
this time. We have spent a lot of time looking for things, more than
ever before.

Weather: More wind today than before, but still warm and pleasant.
Beautiful evening.

Crew Physical Status: Even dirtier, and much more tired.

EVA: John L and Kevin took the rover up the main road to explore. As
Kevin is a new CapCOM, this trip will help him support crew EVA
planning.

Reports to be file: Sol Summary and Engineering Report. Still no crew
photos, but we are determined to take one in the morning.

Support Requested: None at this time.

Operations Report – November 3rd

2017-11-03

Engineering Report – Friday 3 Nov. 2017

Refit Crew: Dr.Shannon Rupert, Kay Radzik Warren, Kevin Seidler, John
Lillquist, and John Stone

******************************************************************
STATUS OF HAB:
Water: Full tank, 500 gallons
Water on trailer: Both tanks empty
Diesel for generator: 75%
Solar Power: 100% SOC
Generator: on from 4 pm.  Last person up tonight will shut it off.
Propane: big tank = 58%, small tank = 40%
Fuel for ATV’s (Jerry Cans): All gas unusable. DO NOT USE FUEL WITH
ETHANOL IN IT
Fuel for Hab Car: Full tank, new battery

******************************************************************
Tasks done today:
1) GreenHab: grow lights installed; shade cloth installed on south
wall; +/- 2 dozen pots planted and watered.
2) Toilet tank flushed out.
3) Tunnel between GreenHab and intersection to Science Lab removed and replaced.
4) Generator jump started, oil was topped off, and it is in working condition.
5) Electrical receptacles checked, and we found that the receptacles
in the toilet, shower room, and alcove in front are not working. Lack
of GFCI outlets in wet areas noted.
6) Replacement lid for vent/skylight fabricated.
7) Porthole in south airlock exterior door had fallen out due to
extreme wind conditions (and the fact that it was just sitting in the
opening and not fastened mechanically, just siliconed), but was
undamaged. So it was put back in with carabiners slotted through holes
on each side so that it won’t fall out again.

******************************************************************
Tasks that need to be done:
A. Water heater is still not working. Per Curtis, it needs to be fully
replaced. Installation expected soon.
B. Need more water in spare tanks.

Sol Summary – November 2nd

Sol Summary Report for sol 0.2

November 2, 2017

Here is when it gets rough!

Shannon Rupert, MDRS Director

Mission Status: We accomplished quite a lot today, considering we are
a small group.  See Kay’s engineering report for details.

Sol Activity Summary: The rest of the crew were up and outside working
when I arrived.  We worked on the tunnel until about 10 am, when a
strange vehicle drove up.  We had a media visit from Voice of America,
and I was their interviewee and there I was, looking much as I usually
do at MDRS, oh so fashionable (stop laughing, all you people who know
me), with my hair a spectacular mess!  I did get to brush my teeth
before filming.  It went very well, in the end, although I normally
try my best to get out of being on camera.  This will be a series of
videos broadcast in Russia.  We continued to work throughout the day,
and although the afternoons are supposed to be free time, this
dedicated group just kept working.  We were going to have pizza for
dinner, but our order did not get to the person who was cooking them,
so we ended up with a bunch of wonderful junk food instead.

Look Ahead Plan:  We need to get water here for the crew who arrive on
Saturday, so that will be our big push tomorrow.

Anomalies in work:  I noticed that this afternoon, with full sun, our
state of charge on the solar system was 53%, which is considerably
lower that it has been.  There was no real extra pull on the power
supply today, so I’m not sure what happened.  With the generator
currently non-functional we are dependent on the solar only.  I am
hoping DG will be able to work on the generator before the crew
arrives.  Also we cannot find the hitch for the HabCar.

Weather:  The weather continues to be unusually warm and pleasant.
This evening the wind has kicked up a little more than yesterday, but
it’s still a nice day.  Almost full moon tonight.

Crew Physical Status:  Dirtier, and a little tired.

EVA: No EVAs

Reports to be file: Sol Summary and Engineering Report. We talked
about getting a crew photo, but then just went back to work.

Support Requested: None at this time.